### v2

# A Chinese restaurant process for multiset permutations

Abstract
Multisets are like sets, except that they can contain multiple copies of their elements. If there are $n_{i}$ copies of $i$, $1\leq i\leq t$, in multiset $M_{t}$, then there are $\binom{n_{1}+\cdots+n_{t}}{n_{1},\ldots,n_{t}}$ possible permutations of $M_{t}$. Knuth showed how to factor any multiset permutation into cycles. For fixed $n_{i}$, $i\geq 1$, we show how to adapt the Chinese restaurant process, which generates random permutations on $n$ elements with weighting $\theta^{\#\,{\rm cycles}}$, $\theta>0$, sequentially for $n=1,2,\ldots$, to the multiset case, where we fix the $n_{i}$ and build permutations on $M_{t}$ sequentially for $t=1,2,\ldots$. The number of cycles of a multiset permutation chosen uniformly at random, i.e. $\theta=1$, has distribution given by the sum of independent negative hypergeometric distributed random variables. For all $\theta>0$, and under the assumption that $n_{i}=O(1)$, we show a central limit theorem as $t\to\infty$ for the number of cycles.
60F05,
keywords:
keywords:
[MSC2020]
volume: 0issue: 0articletype: research-article
[…]
Research Article
1 Introduction
A growth rule which received considerable attention in the literature is the following preferential attachment algorithm on permutations written as the product of cycles known as the Chinese restaurant process (CRP) [1, 6, 9, 14, 17]. For $n\geq 1$, given a permutation of $n-1]:=\left\{1,2,\ldots,n-1\right\}$ which has been constructed at step n1n-1, element $n$ is either appended to the permutation as a new singleton cycle with probability $\frac{\theta}{\theta+n-1}$, or inserted into a random position within the existing permutation with probability equal to $\frac{n-1}{\theta+n-1}$. Let $S_{n}$ be the set of permutations on $[n]$. The permutation obtained at step $n$ has the Ewens distribution, which is the uniform distribution on $n!$ permutations in the case $\theta=1$, and, in general, has probability distribution $\theta^{|\pi|}/\theta_{(n)}$, $\pi\in S_{n}$, where $\theta_{(n)}:=\theta(\theta+1)\cdots(\theta+n-1)$ and $|\pi|$ is the number of cycles in $\pi$.
Let $\Pi_{n}$ be the random permutation of $[n]$ at the $n$th step of the CRP. The distribution of $\Pi_{n}$ is invariant under conjugations. For different orders the permutations are consistent, in the sense that $\Pi_{m}$, for $m<n$, can be derived from $\Pi_{n}$ by removing elements $m+1,\ldots,n$ from their cycles and deleting empty cycles if necessary.
The CRP expresses every permutation $\pi$ of $[n]$ uniquely as the product
$\pi=(x_{11}\ldots x_{1n_{1}}y_{1})(x_{21}\ldots x_{2n_{2}}y_{2})\cdots(x_{t1}\ldots x_{tn_{t}}y_{t}),\quad t\geq 0,$
where the following two conditions are satisfied:
  1. 1.
    $y_{1}<y_{2}<\ldots<y_{t}$
  2. 2.
    $y_{i}<x_{ij}{\rm\ for\ }1\leq j\leq n_{i},1\leq i\leq t$.
We will see in Section 2 that there is a similar unique expression for permutations of multisets.
The number of cycles in $\Pi_{n}$, denoted by $K_{n}$, has probability generating function (p.g.f.)
${\mathbb{E}}z^{K_{n}}=\frac{(\theta z)_{n}}{(\theta)_{n}},$
which corresponds to the distribution of the sum of $n$ independent Bernoulli variables $K_{n}=\sum_{i=1}^{n}I_{i}$, $I_{i}\sim{\rm Bernoulli}(i/(i+\theta-1))$. We have $I_{i}=1$ exactly when the element $i$ starts a new cycle. As $n\to\infty$,
$K_{n}\sim\theta\log n~~{\rm a.s.},~~~\frac{K_{n}-\theta\log n}{\sqrt{\theta\log n}}\stackrel{{\scriptstyle d}}{{\to}}{\rm N}(0,1).$
In this paper, a CRP will be defined for multiset permutations will be derived using the methods in [17]. Multiset permutations are of fundamental interest in combinatorics [16] and random multiset permutations have been studied [7].
2 Multiset permutations and their factorisations into
cycles
There are at least two different ways of defining cycles of multiset permutations [13, 16]. Knuth’s way [13] will be used in this paper. In this section, we summarise the discussion and results on multiset permutation factorisation taken from Section 5.1.2 of [13] that is needed in this paper. In particular, we are going to use Theorem 1 below to factor multiset permutations. A multiset is like a set except that it can have repetitions of identical elements, for example
$M=\{1,1,1,2,2,3,4,4,4,4\},$
where the set $\{1,2,3,4\}$ is given the usual ordering. A permutation of a multiset is an arrangement of its elements in a row, for example
3 1 2 4 4 1 2 4 1 4.3\ 1\ 2\ 4\ 4\ 1\ 2\ 4\ 1\ 4.
We may write this in two-line notation as
$\left(\begin{array}[]{c c c c c c c c c c}1&1&1&2&2&3&4&4&4&4\\ 3&1&2&4&4&1&2&4&1&4\end{array}\right).$ (1)
The intercalation product $\pi_{1}\perp\pi_{2}$ of two multiset permutations $\pi_{1}$ and $\pi_{2}$ is defined in [4, 8] and is obtained by
  1. 1.
    Expressing $\pi_{1}$ and $\pi_{2}$ in the two-line notation.
  2. 2.
    Juxtaposing these two-line notations.
  3. 3.
    Sorting the columns into nondecreasing order of the top line, in such a way that the sorting is stable, in the sense that left-to-right order in the bottom line is preserved when the corresponding top line elements are equal.
For example
3 1 4 1 2 2 4 4 1 4=3 1 2 4 4 1 2 4 1 43\ 1\ 4\ 1\ 2\ $\perp\$2\ 4\ 4\ 1\ 4=3\ 1\ 2\ 4\ 4\ 1\ 2\ 4\ 1\ 4
since
$\left(\begin{array}[]{c c c c c c}1&1&2&3&4\\ 3&1&4&1&2&\end{array}\right)\perp\left(\begin{array}[]{c c c c c}1&2&4&4&4\\ 2&4&4&1&4\end{array}\right)$
$=$ $\left(\begin{array}[]{c c c c c c c c c c}1&1&1&2&2&3&4&4&4&4\\ 3&1&2&4&4&1&2&4&1&4\end{array}\right)$
For possibly repeated elements $x_{1},x_{2},\ldots,x_{m}$, the cycle $(x_{1},\ldots,x_{n})$ stands for the permutation obtained in two line form by sorting the columns of
$\left(\begin{array}[]{c c c c }x_{1}&x_{2}&\ldots&x_{m}\\ x_{2}&x_{3}&\ldots&x_{1}\end{array}\right)$
in a stable manner. For example, we have
[…]
$=$ $\left(\begin{array}[]{c c c c c c c c c c}1&1&1&2&2&3&4&4&4&4\\ 3&1&2&4&4&1&2&4&1&4\end{array}\right)$
and so (1) is a cycle. For these general cycles, $(x_{1}x_{2}\ldots x_{n})$ is not always the same as $(x_{2}\ldots x_{n}x_{1})$.
Knuth [13] proves the following factorisation of multiset permutations into cycles.
Theorem 1 ([13]).
Let the elements of the multiset $M$ be linearly ordered by the relation <<. Every permutation $\pi$ of $M$ has a unique representation as the intercalation
$\pi=(x_{1,1}\ldots x_{1,n_{1}}y_{1})\perp(x_{2,1}\ldots x_{2,n_{2}}y_{2})\perp\cdots\perp(x_{t-1,1}\ldots x_{t-1,n_{t-1}}y_{t-1})$ (6)
for $t\geq 1$, where the following two conditions are satisfied:
  1. 1.
    $y_{1}\leq y_{2}\leq\ldots\leq y_{t-1}$
  2. 2.
    $y_{i}<x_{i,j}{\rm\ for\ }1\leq j\leq n_{i},\,1\leq i\leq t-1$.
For example, the factorisation of (1) corresponding to Theorem 1 is
[…]
3 The CRP for multiset permutations
Given a fixed sequence of positive integers $n_{i}\geq 1$, $i\geq 1$, we will construct a CRP process on permutations of multisets indexed by integers $t\geq 0$. We define $M(n_{1},\ldots,n_{t})$ to be the multiset which contains $n_{i}$ copies of $i$ for $i\in[t]$, $t\geq 1$, and $\emptyset$ for $t=0$. We give the set of elements $[t]$ of $M(n_{1},\ldots,n_{t})$ the usual ordering. The set of permutations of $M(n_{1},\ldots,n_{t})$ is denoted by $S(n_{1},\ldots,n_{t})$ and is of size
$|S(n_{1},\ldots,n_{t})|=\binom{N_{t}}{n_{1},\ldots,n_{t}}$,
where $N_{t}=n_{1}+\cdots+n_{t}$. Note that $S_{n}=S(\underbrace{{1,1,\ldots,1}}_{\text{$n$ times}}).$ The random multiset permutation $\tilde{\Pi}_{t}$ of our process at the $t$th step is an element of $S(n_{1},\ldots,n_{t})$, represented in accordance with (6). At step $t\in[s]$, there are $n_{t}$ copies of $t$ either to be inserted in the existing permutation $\tilde{\Pi}_{t-1}$ or to be put into singleton cycles. (It is not possible to put more than one copy of $t$ in a new cycle because of condition 2 of Theorem 1.) Suppose that $k_{t}\in\{0\}\cup[n_{t}]$ of the $n_{t}$ copies of $t$ start $k_{t}$ new singleton cycles containing $t$, and the $n_{t}-k_{t}$ other copies of $t$ are inserted into the existing permutation. The $n_{t}-k_{t}$ copies of $t$ can be inserted to the left of any of the $y_{i}$ or $x_{ij}$ in (6). A ‘stars and bars’ argument shows that the number of ways of inserting $n_{t}-k_{t}$ copies of $t$ to the left of an element in the multiset $S(n_{1},\ldots,n_{t-1})$ formatted like (6) is
$\binom{N_{t-1}-1+n_{t}-k_{t}}{n_{t}-k_{t}}$, (7)
recalling that
$\binom{-1}{k}=\left\{\begin{array}[]{l l}1&{\rm if\ }k=0;\\ 0&{\rm if\ }k>1.\end{array}\right$.
We now define the multiset CRP. Given $\theta>0$, define
$F_{t}(\theta)=\sum_{k_{t}=0}^{n_{t}}\binom{N_{t-1}-1+n_{t}-k_{t}}{n_{t}-k_{t}}\theta^{k_{t}}$
The special case
$F_{t}(1)=\sum_{k_{t}=0}^{n_{t}}\binom{N_{t-1}-1+n_{t}-k_{t}}{n_{t}-k_{t}}=\binom{N_{t}}{n_{t}}$, (8)
results from a well known identity; see page 161 of [10]. Note that (8) is the total number of ways of extending the permutation from time t1t-1 to time $t$. At time $t\in[s]$, $k_{t}$ new singleton cycles containing $t$ are created and the other $n_{t}-k_{t}$ copies are inserted into the existing permutation in a specified way with probability $\theta^{k_{t}}/F_{t}(\theta)$. If $n_{t}=1$ for all $t$, then $F_{t}(\theta)=\theta+t-1$ and we recover the usual CRP. An induction argument on $t$ shows that all possible $\pi\in S(n_{1},\ldots,n_{t})$ can be constructed in this manner. This can also be seen from the identity
$\binom{N_{t}}{n_{1},\ldots,n_{t}}=\prod_{s=1}^{t}\binom{N_{s}}{n_{s}}$
Next we will find the distribution of $\tilde{\Pi}_{t}$.
Theorem 2.
Given $\theta>0$, for all $\pi\in S(n_{1},\ldots,n_{t})$ the CRP generates $\pi$ at step $t$ with probability equal to $\theta^{|\pi|}/S$, where
$S=\sum_{\pi\in S(n_{1},\ldots,n_{t})}\theta^{|\pi|}=\,\prod_{s=1}^{t}F_{s}(\theta).$ (9)
Proof.
We will use the method in [17]. An arborescence is a directed, rooted tree with edges oriented in agreement with paths from the root to the leaves. Our arborescence has vertex set $V=\bigcup_{s=0}^{t}\Pi(n_{1},\ldots,n_{s})$ and the root is $\emptyset$ (when $s=0$). The leaves of the arborescence are $\Pi(n_{1},\ldots,n_{t})$. Give each leaf $\pi\in\Pi(n_{1},\ldots,n_{t})$ the weight $\theta^{|\pi|}$, where $|\pi|$ is the number of cycles in $\pi$. There is a directed edge from $\pi_{1}\in\Pi(n_{1},\ldots,n_{s-1})$ to $\pi_{2}\in\Pi(n_{1},\ldots,n_{s})$, $s\in[t]$, if and only if, for some $k_{s}\in\{0\}\cup[n_{s}]$, $\pi_{2}$ is obtained from $\pi_{1}$ by inserting $n_{s}-k_{s}$ copies of $s$ to the left of elements of $\pi_{1}$, when $\pi_{1}$ is written as (1), and creating $k_{s}$ singleton cycles each containing $s$. For each $\pi\in V$, define $Q(\pi)$ to be the sum of the weights of all leaves $\pi^{\prime}\in\Pi(n_{1},\ldots,n_{t})$ for which there is a directed path from $\pi$ to $\pi^{\prime}$. For $\pi\in\Pi(n_{1},\ldots,n_{s})$, we have
$Q(\pi)=\theta^{|\pi|}\prod_{u=s+1}^{t}F_{u}(\theta),$ (10)
which follows immediately from the interpretation of the binomial coefficient preceding (7). Taking $\pi=\emptyset$ and $s=0$ in (10) gives (9). With $\pi_{1}$ and $\pi_{2}$ taken as above, we have $|\pi_{2}|=|\pi_{1}|+k_{s}$, and therefore
$p_{\pi_{1},\pi_{2}}:=\frac{Q(\pi_{2})}{Q(\pi_{1})}=\frac{\theta^{k_{s}}}{F_{s}(\theta)}.$ (11)
These are the transition probabilities of the CRP. Theorem 1 of [17] states that if we put transition probabilities (11) on the edges of the arborescence, and let the leaves be absorbing states, then we obtain a Markov chain whose probability of absorption on a given leaf is proportional to the weight of the leaf. By our choice of weight we are done. ∎
For different steps the permutations are consistent, in the sense that $\tilde{\Pi}_{s}$, for $s<t$, can be derived from $\tilde{\Pi}_{t}$ by removing elements $s+1,\ldots,t$ from their cycles and deleting empty cycles if necessary.
4 The distribution of the number of cycles
Let $X_{t}$ denote the number of new cycles added in moving from time t1t-1 to time $t$, $t\geq 1$, which is the choice of $k_{t}$ above. The $X_{t}$ are independent and have p.g.f.
${\rm I\mkern-3.5muP}(X_{t}=k_{t})=\frac{\theta^{k_{t}}}{F_{t}(\theta)}\binom{N_{t-1}-1+n_{t}-k_{t}}{n_{t}-k_{t}},\quad k_{t}\in\{0\}\cup[n_{t}].$ (12)
If $\theta=1$, then, by (8), we have
${\rm I\mkern-3.5muP}(X_{t}=k_{t})=\frac{1}{\binom{N_{t}}{n_{t}}}\binom{N_{t-1}-1+n_{t}-k_{t}}{n_{t}-k_{t}}=\frac{\binom{n_{t}}{k_{t}}}{\binom{N_{t}}{k_{t}}}\frac{N_{t}-n_{t}}{N_{t}-k_{t}}$ (13)
[…]
${\mathbb{P}}(Y=\kappa)=\frac{\binom{M}{r-1}\binom{N-M}{\kappa-r}}{\binom{N}{\kappa-1}}\cdot\frac{M-r+1}{N-\kappa+1}$ (14)
It is easily checked that (13) is (14) with $N=N_{t}$< $M=N_{t}-n_{t}=N_{t-1}$, $r=1$ and $\kappa=k+1$. The expectation of $Y$ is
${\mathbb{E}}(Y)=r\frac{N+1}{M+1},$
the variance of $Y$ is
${\rm Var}(Y)=\frac{r(N-M)(N+1)(M+1-r)}{(M+1)^{2}(M+2)}$
and the third centred moment (see [11, 12]) is
${\mathbb{E}}([Y-{\mathbb{E}}(Y)]^{3})={\rm Var}(Y)\cdot\frac{(2N-M+1)(M+1-2r)}{(M+1)(M+3)}.$
Therefore,
${\mathbb{E}}(X_{t})={\mathbb{E}}(Y_{t})-1=\frac{N_{t}+1}{N_{t-1}+1}-1=\frac{n_{t}}{N_{t-1}+1},$ (15)
${\rm Var}(X_{t})={\rm Var}(Y_{t})=\frac{n_{t}(N_{t}+1)N_{t-1}}{(N_{t-1}+1)^{2}(N_{t-1}+2)}$ (16)
and
${\mathbb{E}}([X_{t}-{\mathbb{E}}(X_{t})]^{3})$ $=$ ${\mathbb{E}}([Y_{t}-{\mathbb{E}}(Y_{t})]^{3})$ (17)
$=$ ${\rm Var}(Y_{t})\cdot\frac{(n_{t}+N_{t}+1)(N_{t-1}-1)}{(N_{t-1}+1)(N_{t-1}+3)}.$
It follows that the number of cycles $K_{t}$ has expectation
${\mathbb{E}}(K_{t})=\sum_{s=1}^{s}\frac{n_{s}}{N_{s-1}+1}$
and variance
${\rm Var}(K_{t})=\sum_{s=1}^{t}\frac{n_{s}(N_{s}+1)N_{s-1}}{(N_{s-1}+1)^{2}(N_{s-1}+2)}.$
Note that $n_{t}$ only appears in the numerators in (15) and (16), and therefore, by increasing $n_{t}$, with $n_{1},n_{2},\ldots,n_{t-1}$ fixed, we can make the expectaion and variance as large as desired.
Given sequences $a_{t}$ and $b_{t}$, we use the notation $a_{t}=O(b_{t})$ to mean $a_{t}\leq Cb_{t}$ for a constant $C>0$; $a_{t}\asymp b_{t}$ to mean $C_{1}b_{t}\leq a_{t}\leq C_{2}b_{t}$ for constants $C_{1}>0$, $C_{2}>0$; and $a_{t}\sim b_{t}$ to mean $\lim_{t\to\infty}a_{t}/b_{t}=1$. We will put the bound the $n_{t}=O(1)$ to give a central limit theorem for $K_{s}$ for all $\theta$.
Theorem 3.
If $\theta=1$ and $n_{i}=n\geq 1$ for all $i\geq 1$, then
${\mathbb{E}}(K_{t})\sim\log t\quad{\rm Var}(K_{t})\sim\log t.$ (18)
For general $\theta$, let $n_{1},n_{2},\ldots,$ be given, $n_{i}\geq 1$ for all $i$, and suppose
$n_{i}=O(1).$ (19)
Then,
${\mathbb{E}}(K_{t})\asymp\log t\quad{\rm Var}(K_{t})\asymp\log t,$ (20)
and
$K_{n}\sim{\mathbb{E}}(K_{n})~~{\rm a.s.}~~~~~~\frac{K_{t}-{\mathbb{E}}(K_{t})}{\sqrt{\rm Var}(K_{t})}\stackrel{{\scriptstyle d}}{{\to}}{\rm N}(0,1).$ (21)
Proof.
First we prove the theorem for $\theta=1$. From (15), (16) and (19) we get ${\mathbb{E}}(X_{t})\asymp 1/t$ and ${\rm Var}(X_{t})\asymp 1/t$, which result in (20), after which the Lindeberg-Feller Theorem easily shows the central limit theorem in (21), because $X_{t}\leq n_{t}=O(1)$. If we impose the condition $n_{i}=n$, for $i\geq 1$, then we immediately obtain (18). For $\theta$ general, we reduce to the case $\theta=1$. This follows from noting that in (12), $\theta^{k_{t}}\asymp 1$ and $F_{t}(\theta)\asymp F_{t}(1)$ uniformly over $t\geq 1$, by (19), and so (20) still holds. For the almost sure convergence, modify the proof of Theorem 2.3.8 of [5]. ∎
The following theorem can get a Central Limit Theorem when $n_{t}$ is unbounded.
Theorem 4.
Suppose $\theta=1$, $n_{t}=O(N_{t-1})$, and $\sum_{t=1}^{\infty}\frac{n_{t}}{N_{t-1}}=\infty$. Then, (21) holds.
Proof.
In order to prove limiting normality, we will verify Lyapanov’s Condition in [3] with $\delta=1$, by showing that
$\lim_{t\to\infty}\,\frac{1}{{\rm Var}(K_{t})^{3/2}}\sum_{s=1}^{t}{\mathbb{E}}([X_{s}-{\mathbb{E}}(X_{s})]^{3})=0.$ (22)
The assumption $n_{t}=O(N_{t-1})$, (16) and (17) produces
${\mathbb{E}}([X_{t}-{\mathbb{E}}(X_{t})]^{3}\asymp{\rm Var}(X_{t})\asymp n_{t}/N_{t-1}.$
Hence,
${\rm Var}(K_{t})\asymp\sum_{s=1}^{t}{\mathbb{E}}([X_{s}-{\mathbb{E}}(X_{s})]^{3})\asymp\sum_{s=1}^{t}\frac{n_{s}}{N_{s-1}}$
and (22) holds by the second assumption. ∎
An example where Theorem 4 holds is obtained by setting $\theta=1$, $n_{t}=\lfloor Ct^{\alpha}\rfloor$, for constants $C>0$, $\alpha>0$, where $\lfloor x\rfloor$ is the smallest integer greater or equal to $x$. Then, $n_{t}=Ct^{\alpha}+O(1)$ and $N_{t}=Ct^{\alpha+1}/(\alpha+1)+O(t^{\alpha})$ and the assumptions of Theorem 4 hold. The expectation and variance are asymptotically
${\mathbb{E}}(K_{t})\sim(\alpha+1)\log t\quad{\rm Var}(K_{t})\sim(\alpha+1)\log t,$
5 Discussion
Potentially, improvements might be made on these results. The CRP we have defined places all $n_{t}$ elements labelled $t$ at the same time. Placing them sequentially seems like a difficult problem. It would be interesting to get better estimates for ${\mathbb{E}}(K_{t})$ and ${\rm Var}(K_{t})$, especially when $\theta\neq 1$. Perhaps the hypotheses of Theorem 4 could be weakened.
In the case of random permutations, there is a Poisson approximation of $K_{n}$. The total variation distance between the law ${\cal L}(K_{n})$ of $K_{n}$ and the ${\rm Poisson}({\mathbb{E}}(K_{n}))$ distribution is of order
$d_{\rm TV}({\cal L}(K_{n}),{\rm Poisson}({\mathbb{E}}(K_{n})))$ $:=$ $\frac{1}{2}\sum_{k=1}^{n}\left|{\mathbb{P}}(K_{n}=k)-\exp\left(-{\mathbb{E}}(K_{n})\right)\frac{{\mathbb{E}}(K_{n})^{k}}{k!}\right|$
$\asymp$ $\frac{1}{\log n};$
see [2]. By considering the process of cycle counts $(C_{1}(n),C_{2}(n),\ldots,C_{n}(n))$, where $C_{i}(n)$ is the number of cycles of size $i$, we may write $K_{n}=\sum_{i=1}^{n}C_{i}(n)$, and, moreover, one may show that
$\frac{\sum_{i=1}^{\lfloor n^{t}\rfloor}C_{i}(n)-t\log n}{\sqrt{\log n}}$
is asymptotically standard Brownian motion; see [1]. If suitable approximations could be made to the component count process of multiset permutations, then similar progress might be made for the results obtained in this paper when $\theta=1$.

### v3

@@
 1 Introduction
-A growth rule which received considerable attention in the literature is the following preferential attachment algorithm on permutations written as the product of cycles known as the Chinese restaurant process (CRP) [1, 6, 9, 14, 17].
+A growth rule which received considerable attention in the literature is the following preferential attachment algorithm on permutations written as the product of cycles known as the Chinese restaurant process (CRP) [ABT, BMMU, Feng, GI, GW, GS, IKMW, CSP, Stark].
 For $n\geq 1$, given a permutation of $n-1]:=\left\{1,2,\ldots,n-1\right\}$ which has been constructed at step n1n-1, element $n$ is either appended to the permutation as a new singleton cycle with probability $\frac{\theta}{\theta+n-1}$, or inserted into a random position within the existing permutation with probability equal to $\frac{n-1}{\theta+n-1}$.
@@
 Let $\Pi_{n}$ be the random permutation of $[n]$ at the $n$th step of the CRP.
-The distribution of $\Pi_{n}$ is invariant under conjugations.
+The distribution of $\Pi_{n}$ is invariant under relabelling.
 For different orders the permutations are consistent, in the sense that $\Pi_{m}$, for $m<n$, can be derived from $\Pi_{n}$ by removing elements $m+1,\ldots,n$ from their cycles and deleting empty cycles if necessary.
 The CRP expresses every permutation $\pi$ of $[n]$ uniquely as the product
-$\pi=(x_{11}\ldots x_{1n_{1}}y_{1})(x_{21}\ldots x_{2n_{2}}y_{2})\cdots(x_{t1}\ldots x_{tn_{t}}y_{t}),\quad t\geq 0,$
+$\pi=(x_{1,1}\ldots x_{1,m_{1}}y_{1})(x_{2,1}\ldots x_{2,m_{2}}y_{2})\cdots(x_{\tau,1}\ldots x_{\tau,m_{\tau}}y_{\tau}),\quad\tau\geq 1,$
 where the following two conditions are satisfied:
   1. 1.
-    $y_{1}<y_{2}<\ldots<y_{t}$
+    $y_{1}<y_{2}<\ldots<y_{\tau}$
   2. 2.
-    $y_{i}<x_{ij}{\rm\ for\ }1\leq j\leq n_{i},1\leq i\leq t$.
+    $y_{i}<x_{ij}{\rm\ for\ }1\leq j\leq m_{i},1\leq i\leq\tau$.
 We will see in Section 2 that there is a similar unique expression for permutations of multisets.
@@
 ${\mathbb{E}}z^{K_{n}}=\frac{(\theta z)_{n}}{(\theta)_{n}},$
-which corresponds to the distribution of the sum of $n$ independent Bernoulli variables $K_{n}=\sum_{i=1}^{n}I_{i}$, $I_{i}\sim{\rm Bernoulli}(i/(i+\theta-1))$.
+which corresponds to the distribution of the sum of $n$ independent Bernoulli variables $K_{n}=\sum_{i=1}^{n}I_{i}$, $I_{i}\sim{\rm Bernoulli}(\theta/(i+\theta-1))$; see [ABT].
 We have $I_{i}=1$ exactly when the element $i$ starts a new cycle.
 As $n\to\infty$,
-$K_{n}\sim\theta\log n~~{\rm a.s.},~~~\frac{K_{n}-\theta\log n}{\sqrt{\theta\log n}}\stackrel{{\scriptstyle d}}{{\to}}{\rm N}(0,1).$
-In this paper, a CRP will be defined for multiset permutations will be derived using the methods in [17].
-Multiset permutations are of fundamental interest in combinatorics [16] and random multiset permutations have been studied [7].
+$K_{n}\sim\theta\log n~~{\rm a.s.},~~~\frac{K_{n}-\theta\log n}{\sqrt{\theta\log n}}\stackrel{{\scriptstyle d}}{{\to}}{\rm N}(0,1),$
+where $\stackrel{{\scriptstyle d}}{{\to}}$ denotes convergence in distribution.
+In this paper, a CRP will be defined for multiset permutations and derived using the methods in [Stark].
+As motivation, we note that multiset permutations are of fundamental interest in combinatorics [Stanley] and random multiset permutations have been studied previously [Feray].
 2 Multiset permutations and their factorisations into
 cycles
-There are at least two different ways of defining cycles of multiset permutations [13, 16].
-Knuth’s way [13] will be used in this paper.
-In this section, we summarise the discussion and results on multiset permutation factorisation taken from Section 5.1.2 of [13] that is needed in this paper.
+There are at least two different ways of defining cycles of multiset permutations [Knuth, Stanley].
+Knuth’s way [Knuth] will be used in this paper.
+In this section, we summarise the discussion and results on multiset permutation factorisation taken from Section 5.1.2 of [Knuth] that is needed in this paper.
 In particular, we are going to use Theorem 1 below to factor multiset permutations.
@@
 $\left(\begin{array}[]{c c c c c c c c c c}1&1&1&2&2&3&4&4&4&4\\ 3&1&2&4&4&1&2&4&1&4\end{array}\right).$ (1)
-The intercalation product $\pi_{1}\perp\pi_{2}$ of two multiset permutations $\pi_{1}$ and $\pi_{2}$ is defined in [4, 8] and is obtained by
+The intercalation product $\pi_{1}\perp\pi_{2}$ of two multiset permutations $\pi_{1}$ and $\pi_{2}$ is defined in [CF, Foata] and is obtained by
   1. 1.
@@
     Sorting the columns into nondecreasing order of the top line, in such a way that the sorting is stable, in the sense that left-to-right order in the bottom line is preserved when the corresponding top line elements are equal.
-For example
-3 1 4 1 2 2 4 4 1 4=3 1 2 4 4 1 2 4 1 43\ 1\ 4\ 1\ 2\ $\perp\$2\ 4\ 4\ 1\ 4=3\ 1\ 2\ 4\ 4\ 1\ 2\ 4\ 1\ 4
-since
-$\left(\begin{array}[]{c c c c c c}1&1&2&3&4\\ 3&1&4&1&2&\end{array}\right)\perp\left(\begin{array}[]{c c c c c}1&2&4&4&4\\ 2&4&4&1&4\end{array}\right)$
+For example,
+$\left(\begin{array}[]{c c c c c c}1&1&2&3&4\\ 3&1&4&1&2&\end{array}\right)\top\left(\begin{array}[]{c c c c c}1&2&4&4&4\\ 2&4&4&1&4\end{array}\right)$
 $=$ $\left(\begin{array}[]{c c c c c c c c c c}1&1&1&2&2&3&4&4&4&4\\ 3&1&2&4&4&1&2&4&1&4\end{array}\right)$
@@
 […]
-$=$ $\left(\begin{array}[]{c c c c c c c c c c}1&1&1&2&2&3&4&4&4&4\\ 3&1&2&4&4&1&2&4&1&4\end{array}\right)$
-and so (1) is a cycle. For these general cycles, $(x_{1}x_{2}\ldots x_{n})$ is not always the same as $(x_{2}\ldots x_{n}x_{1})$.
-Knuth [13] proves the following factorisation of multiset permutations into cycles.
-Theorem 1 ([13]).
+$=$ $\left(\begin{array}[]{c c c c c c c c c c}4&2&4&4&1&3&1&1&2&4\\ 2&4&4&1&3&1&1&2&4&4\end{array}\right)$ (7)
+$=$ $\left(\begin{array}[]{c c c c c c c c c c}1&1&1&2&2&3&4&4&4&4\\ 3&1&2&4&4&1&2&4&1&4\end{array}\right)$ (10)
+and so (1) is a cycle.
+For these general cycles, $(x_{1}x_{2}\ldots x_{n})$ is not always the same as $(x_{2}\ldots x_{n}x_{1})$ and it can happen that a cycle can be written as the intercalation product of other cycles, as we shall see at the end of this section.
+Knuth [Knuth] proves the following factorisation of multiset permutations into cycles.
+Theorem 1 ([Knuth]).
 Let the elements of the multiset $M$ be linearly ordered by the relation <<. Every permutation $\pi$ of $M$ has a unique representation as the intercalation
-$\pi=(x_{1,1}\ldots x_{1,n_{1}}y_{1})\perp(x_{2,1}\ldots x_{2,n_{2}}y_{2})\perp\cdots\perp(x_{t-1,1}\ldots x_{t-1,n_{t-1}}y_{t-1})$ (6)
-for $t\geq 1$, where the following two conditions are satisfied:
+$\pi=(x_{1,1}\ldots x_{1,m_{1}}y_{1})\top(x_{2,1}\ldots x_{2,m_{2}}y_{2})\top\cdots\top(x_{\tau,1}\ldots x_{\tau,n_{\tau}}y_{\tau}),\quad\tau\geq 1,$ (11)
+where the following two conditions are satisfied:
   1. 1.
-    $y_{1}\leq y_{2}\leq\ldots\leq y_{t-1}$
+    $y_{1}\leq y_{2}\leq\ldots\leq y_{\tau}$
   2. 2.
-    $y_{i}<x_{i,j}{\rm\ for\ }1\leq j\leq n_{i},\,1\leq i\leq t-1$.
-For example, the factorisation of (1) corresponding to Theorem 1 is
-[…]
+    $y_{i}<x_{i,j}{\rm\ for\ }1\leq j\leq m_{i},\,1\leq i\leq\tau$.
+Note that (4) is not in the form (11). The factorisation of (1) corresponding to Theorem 1 is
+$(3\ 1)\top(1)\top(2\ 4\ 2\ 4\ 4\ 1)\top(4)$
+$=$ $\left(\begin{array}[]{c c}3&1\\ 1&3\end{array}\right)\top\left(\begin{array}[]{c}1\\ 1\end{array}\right)\top\left(\begin{array}[]{c c c c c c}2&4&2&4&4&1\\ 4&2&4&4&1&2\end{array}\right)\top\left(\begin{array}[]{c}4\\ 4\end{array}\right)$
+$=$ $\left(\begin{array}[]{c c c c c c c c c c}3&1&1&2&4&2&4&4&1&4\\ 1&3&1&4&2&4&4&1&2&4\end{array}\right)$
+$=$ $\left(\begin{array}[]{c c c c c c c c c c}1&1&1&2&2&3&4&4&4&4\\ 3&1&2&4&4&1&2&4&1&4\end{array}\right).$
 3 The CRP for multiset permutations
@@
 Note that $S_{n}=S(\underbrace{{1,1,\ldots,1}}_{\text{$n$ times}}).$ The random multiset permutation $\tilde{\Pi}_{t}$ of our process at the $t$th step is an element of $S(n_{1},\ldots,n_{t})$, represented in accordance with (11).
-At step $t\in[s]$, there are $n_{t}$ copies of $t$ either to be inserted in the existing permutation $\tilde{\Pi}_{t-1}$ or to be put into singleton cycles.
+At step $t\geq 1$, there are $n_{t}$ copies of $t$ either to be inserted in the existing permutation $\tilde{\Pi}_{t-1}$ or to be put into singleton cycles.
 (It is not possible to put more than one copy of $t$ in a new cycle because of condition 2 of Theorem 1.) Suppose that $k_{t}\in\{0\}\cup[n_{t}]$ of the $n_{t}$ copies of $t$ start $k_{t}$ new singleton cycles containing $t$, and the $n_{t}-k_{t}$ other copies of $t$ are inserted into the existing permutation.
@@
 The special case
-$F_{t}(1)=\sum_{k_{t}=0}^{n_{t}}\binom{N_{t-1}-1+n_{t}-k_{t}}{n_{t}-k_{t}}=\binom{N_{t}}{n_{t}}$, (8)
-results from a well known identity; see page 161 of [10].
+$F_{t}(1)=\sum_{k_{t}=0}^{n_{t}}\binom{N_{t-1}-1+n_{t}-k_{t}}{n_{t}-k_{t}}=\binom{N_{t}}{n_{t}}$. (16)
+results from the well known identity
+$\sum_{\rho=0}^{\nu}\binom{\mu+\rho}{\rho}=\binom{\mu+\nu+1}{\nu},\quad\nu,\mu\geq 0\ {\rm integers};$ (17)
+see page 161 of [GKP].
 Note that (16) is the total number of ways of extending the permutation from time t1t-1 to time $t$.
-At time $t\in[s]$, $k_{t}$ new singleton cycles containing $t$ are created and the other $n_{t}-k_{t}$ copies are inserted into the existing permutation in a specified way with probability $\theta^{k_{t}}/F_{t}(\theta)$.
+At time $t\geq 1$, $k_{t}$ new singleton cycles containing $t$ are created and the other $n_{t}-k_{t}$ copies are inserted into the existing permutation in a specified way with probability $\theta^{k_{t}}/F_{t}(\theta)$.
 If $n_{t}=1$ for all $t$, then $F_{t}(\theta)=\theta+t-1$ and we recover the usual CRP.
@@
 Proof.
-We will use the method in [17].
+We will use the method in [Stark].
 An arborescence is a directed, rooted tree with edges oriented in agreement with paths from the root to the leaves.
@@
 These are the transition probabilities of the CRP.
-Theorem 1 of [17] states that if we put transition probabilities (11) on the edges of the arborescence, and let the leaves be absorbing states, then we obtain a Markov chain whose probability of absorption on a given leaf is proportional to the weight of the leaf.
+Theorem 1 of [Stark] states that if we put transition probabilities (20) on the edges of the arborescence, and let the leaves be absorbing states, then we obtain a Markov chain whose probability of absorption on a given leaf is proportional to the weight of the leaf.
 By our choice of weight we are done.
@@
 4 The distribution of the number of cycles
-Let $X_{t}$ denote the number of new cycles added in moving from time t1t-1 to time $t$, $t\geq 1$, which is the choice of $k_{t}$ above. The $X_{t}$ are independent and have p.g.f.
+Let $X_{t}$ denote the number of new cycles added in moving from time t1t-1 to time $t$, $t\geq 1$, which is the choice of $k_{t}$ above. The $X_{t}$ are independent with distributions given by
 ${\rm I\mkern-3.5muP}(X_{t}=k_{t})=\frac{\theta^{k_{t}}}{F_{t}(\theta)}\binom{N_{t-1}-1+n_{t}-k_{t}}{n_{t}-k_{t}},\quad k_{t}\in\{0\}\cup[n_{t}].$ (21)
@@
 ${\mathbb{P}}(Y=\kappa)=\frac{\binom{M}{r-1}\binom{N-M}{\kappa-r}}{\binom{N}{\kappa-1}}\cdot\frac{M-r+1}{N-\kappa+1}$ (23)
-It is easily checked that (13) is (14) with $N=N_{t}$< $M=N_{t}-n_{t}=N_{t-1}$, $r=1$ and $\kappa=k+1$. The expectation of $Y$ is
+It is easily checked that (22) is (23) with $N=N_{t}$, $M=N_{t}-n_{t}=N_{t-1}$, $r=1$ and $\kappa=k+1$. The expectation of $Y$ is
 ${\mathbb{E}}(Y)=r\frac{N+1}{M+1},$
@@
 ${\rm Var}(Y)=\frac{r(N-M)(N+1)(M+1-r)}{(M+1)^{2}(M+2)}$
-and the third centred moment (see [11, 12]) is
+and the third centred moment (see [JKK, KK]) is
 ${\mathbb{E}}([Y-{\mathbb{E}}(Y)]^{3})={\rm Var}(Y)\cdot\frac{(2N-M+1)(M+1-2r)}{(M+1)(M+3)}.$
@@
 $=$ ${\rm Var}(Y_{t})\cdot\frac{(n_{t}+N_{t}+1)(N_{t-1}-1)}{(N_{t-1}+1)(N_{t-1}+3)}.$
-It follows that the number of cycles $K_{t}$ has expectation
-${\mathbb{E}}(K_{t})=\sum_{s=1}^{s}\frac{n_{s}}{N_{s-1}+1}$
+It follows that the number of cycles $K_{t}=\sum_{s=1}^{t}X_{t}$ has expectation
+${\mathbb{E}}(K_{t})=\sum_{s=1}^{t}\frac{n_{s}}{N_{s-1}+1}$
 and variance
 ${\rm Var}(K_{t})=\sum_{s=1}^{t}\frac{n_{s}(N_{s}+1)N_{s-1}}{(N_{s-1}+1)^{2}(N_{s-1}+2)}.$
-Note that $n_{t}$ only appears in the numerators in (15) and (16), and therefore, by increasing $n_{t}$, with $n_{1},n_{2},\ldots,n_{t-1}$ fixed, we can make the expectaion and variance as large as desired.
 Given sequences $a_{t}$ and $b_{t}$, we use the notation $a_{t}=O(b_{t})$ to mean $a_{t}\leq Cb_{t}$ for a constant $C>0$; $a_{t}\asymp b_{t}$ to mean $C_{1}b_{t}\leq a_{t}\leq C_{2}b_{t}$ for constants $C_{1}>0$, $C_{2}>0$; and $a_{t}\sim b_{t}$ to mean $\lim_{t\to\infty}a_{t}/b_{t}=1$.
-We will put the bound the $n_{t}=O(1)$ to give a central limit theorem for $K_{s}$ for all $\theta$.
+We also define $a_{t}=o(b_{t})$ to mean $\lim_{t\to\infty}a_{t}/b_{t}=0$.
+We will put the bound the $n_{t}=O(1)$ to give a central limit theorem for $K_{t}$ for all $\theta$.
 Theorem 3.
 If $\theta=1$ and $n_{i}=n\geq 1$ for all $i\geq 1$, then
-${\mathbb{E}}(K_{t})\sim\log t\quad{\rm Var}(K_{t})\sim\log t.$ (18)
+${\mathbb{E}}(K_{t})\sim\log t,\quad{\rm Var}(K_{t})\sim\log t.$ (27)
 For general $\theta$, let $n_{1},n_{2},\ldots,$ be given, $n_{i}\geq 1$ for all $i$, and suppose
@@
 Then,
-${\mathbb{E}}(K_{t})\asymp\log t\quad{\rm Var}(K_{t})\asymp\log t,$ (20)
+${\mathbb{E}}(K_{t})\asymp\log t,\quad{\rm Var}(K_{t})\asymp\log t,$ (29)
 and
-$K_{n}\sim{\mathbb{E}}(K_{n})~~{\rm a.s.}~~~~~~\frac{K_{t}-{\mathbb{E}}(K_{t})}{\sqrt{\rm Var}(K_{t})}\stackrel{{\scriptstyle d}}{{\to}}{\rm N}(0,1).$ (21)
+$K_{n}\sim{\mathbb{E}}(K_{n})~~{\rm a.s.},~~~~~~\frac{K_{t}-{\mathbb{E}}(K_{t})}{\sqrt{{\rm Var}(K_{t})}}\stackrel{{\scriptstyle d}}{{\to}}{\rm N}(0,1).$ (30)
 Proof.
@@
 This follows from noting that in (21), $\theta^{k_{t}}\asymp 1$ and $F_{t}(\theta)\asymp F_{t}(1)$ uniformly over $t\geq 1$, by (28), and so (29) still holds.
-For the almost sure convergence, modify the proof of Theorem 2.3.8 of [5].
+For the almost sure convergence, modify the proof of (6.6) Theorem on page 44 of [durrett].
 ∎
-The following theorem can get a Central Limit Theorem when $n_{t}$ is unbounded.
+By the following theorem a Central Limit Theorem can be obtained even when $n_{t}$ is unbounded, which is not the case in Theorem 3.
 Theorem 4.
-Suppose $\theta=1$, $n_{t}=O(N_{t-1})$, and $\sum_{t=1}^{\infty}\frac{n_{t}}{N_{t-1}}=\infty$. Then, (21) holds.
+Suppose $\theta=1$, $n_{t}=O(N_{t-1})$. Then, (30) holds.
 Proof.
-In order to prove limiting normality, we will verify Lyapanov’s Condition in [3] with $\delta=1$, by showing that
+In order to prove limiting normality, we will verify Lyapounov’s Condition in [Bill] with $\delta=1$, by showing that
 $\lim_{t\to\infty}\,\frac{1}{{\rm Var}(K_{t})^{3/2}}\sum_{s=1}^{t}{\mathbb{E}}([X_{s}-{\mathbb{E}}(X_{s})]^{3})=0.$ (31)
 The assumption $n_{t}=O(N_{t-1})$, (25) and (26) produces
-${\mathbb{E}}([X_{t}-{\mathbb{E}}(X_{t})]^{3}\asymp{\rm Var}(X_{t})\asymp n_{t}/N_{t-1}.$
-Hence,
-${\rm Var}(K_{t})\asymp\sum_{s=1}^{t}{\mathbb{E}}([X_{s}-{\mathbb{E}}(X_{s})]^{3})\asymp\sum_{s=1}^{t}\frac{n_{s}}{N_{s-1}}$
-and (22) holds by the second assumption. ∎
-An example where Theorem 4 holds is obtained by setting $\theta=1$, $n_{t}=\lfloor Ct^{\alpha}\rfloor$, for constants $C>0$, $\alpha>0$, where $\lfloor x\rfloor$ is the smallest integer greater or equal to $x$.
+${\mathbb{E}}([X_{t}-{\mathbb{E}}(X_{t})]^{3}\asymp{\rm Var}(X_{t})\asymp n_{t}/N_{t-1},\quad t\geq 2.$
+We have the lower bound
+$\sum_{s=2}^{t}\frac{n_{s}}{N_{s-1}}$ $\geq$ $\sum_{s=2}^{t}\ln\left(1+\frac{n_{s}}{N_{s-1}}\right)$
+$=$ $\sum_{s=2}^{t}\left(\ln(N_{s})-\ln(N_{s-1}\right)$
+$=$ $\ln(N_{t})-\ln(N_{1})\to\infty,$
+where the convergence to infinity follows from the assumption $n_{t}\geq 1$ for all $t\geq 1$. Hence,
+${\rm Var}(K_{t})\asymp\sum_{s=1}^{t}{\mathbb{E}}([X_{s}-{\mathbb{E}}(X_{s})]^{3})\asymp\sum_{s=2}^{t}\frac{n_{s}}{N_{s-1}}\to\infty$
+and (31) holds.
+For the almost sure convergence, modify the proof of (6.6) Theorem on page 44 of [durrett] as in the proof of Theorem 3. ∎
+An example where Theorem 4 holds is obtained by setting $\theta=1$, $n_{t}=\lceil Ct^{\alpha}\rceil$, for constants $C>0$, $\alpha>0$, where $\lceil x\rceil$ is the smallest integer greater or equal to $x$.
 Then, $n_{t}=Ct^{\alpha}+O(1)$ and $N_{t}=Ct^{\alpha+1}/(\alpha+1)+O(t^{\alpha})$ and the assumptions of Theorem 4 hold.
 The expectation and variance are asymptotically
-${\mathbb{E}}(K_{t})\sim(\alpha+1)\log t\quad{\rm Var}(K_{t})\sim(\alpha+1)\log t,$
+${\mathbb{E}}(K_{t})\sim(\alpha+1)\log t,\quad{\rm Var}(K_{t})\sim(\alpha+1)\log t.$
+Another example for $\theta=1$ is obtained through regular variation [BGT]. A positive, measurable function $\ell(x)$, defined on $\eta,\infty)$ for $\eta$ real, is slowly varying if
+$\lim_{x\to\infty}\frac{\ell(\lambda x)}{\ell(x)}=1,\quad\forall\lambda>0.$ (32)
+The Uniform Convergence Theorem [BGT] provides that the convergence in (32) is uniform on compact $\lambda$-sets in $(0,\infty)$.
+A positive, measurable function defined on $\eta,\infty)$, is regularly varying with real index $\alpha$ if
+$\lim_{x\to\infty}\frac{f(\lambda x)}{f(x)}=\lambda^{\alpha},\quad\forall\lambda>0.$
+A regularly varying function with index $\alpha$ can be written as
+$f(x)=x^{\alpha}\ell(x),$
+where $\ell(x)$ is a slowly varying function. Suppose that $\ell(x)$ is locally bounded in $\eta,\infty)$, i.e. bounded on compact subsets of $\eta,\infty)$, and $\alpha>-1$. Karamata’s theorem gives
+$\int_{\eta}^{t}x^{\alpha}\ell(x)\,dx\sim t^{\alpha+1}\ell(t)/(\alpha+1),\quad t\to\infty.$
+We further suppose that $\eta\leq 1$, $\alpha\geq 0$, with $\lim_{x\to\infty}\ell(x)=\infty$ in the case $\alpha=0$, and define
+$n_{t}=\Big\lceil\int_{t}^{t+1}x^{\alpha}\ell(x)\,dx\Big\rceil,\quad t\geq 1.$
+The Uniform Convergence Theorem implies that
+$\sup_{t\leq x\leq t+1}\left|\frac{\ell(x)}{\ell(t)}-1\right|\leq\sup_{1\leq\lambda\leq 2}\left|\frac{\ell(\lambda t)}{\ell(t)}-1\right|=o(1),$
+from which
+$\int_{t}^{t+1}x^{\alpha}\ell(x)\,dx$ $=$ $(1+o(1))\ell(t)\int_{t}^{t+1}x^{\alpha}\,dx$
+$=$ $(1+o(1))\ell(t)(t^{\alpha}+o(t^{\alpha}))$
+and
+$n_{t}\sim t^{\alpha}\ell(t).$
+Moreover, Karamata’s theorem gives us
+$N_{t-1}\sim t^{\alpha+1}\ell(t)/(\alpha+1).$
+Hence, the conditions of Theorem 4 hold.
+We will now show that $n_{t}$ may be chosen in such a way that the conclusion of the central limit theorem does not hold for $K_{t}$. Suppose that $\theta=1$ and the $n_{t}$ are chosen such that
+$N_{t-1}=o(n_{t}).$ (33)
+By (24) and (25), we have
+${\mathbb{E}}(X_{t})\sim\frac{n_{t}}{N_{t-1}},\quad{\rm Var}(X_{t})\sim\left(\frac{n_{t}}{N_{t-1}}\right)^{2}.$
+We further choose the $n_{t}$ such that
+$\sum_{s=1}^{t-1}{\rm Var}(X_{s})=o\left({\rm Var}(X_{t})\right),$ (34)
+A particular example of such a sequence is
+$n_{t}=\lceil e^{t^{3}}\rceil=e^{t^{3}}+O(1),$
+from which
+$N_{t}$ $=$ $\sum_{s=1}^{t}e^{s^{3}}+O(t)$
+$=$ $e^{t^{3}}\left(1+\sum_{s=1}^{t-1}e^{s^{3}-t^{3}}\right)+O(t)$
+$=$ $e^{t^{3}}\left(1+O\left(te^{(t-1)^{3}-t^{3}}\right)\right)+O(t)$
+$=$ $e^{t^{3}}(1+o(1)).$
+Clearly, (33) holds and we also have
+$\sum_{s=1}^{t-1}{\rm Var}(X_{s})$ $=$ $O\left(\sum_{s=1}^{t-1}e^{2s^{3}-2(s-1)^{3}}\right)$
+$=$ $O\left(\sum_{s=1}^{t-1}e^{6s^{2}-6s}\right)$
+$=$ $O\left(te^{6(t-1)^{2}-6(t-1)}\right)$
+$=$ $o\left(e^{6t^{2}-6t}\right)$
+and so (34) holds, as well, because ${\rm Var}(X_{t})\sim e^{6t^{2}-6t+2}$.
+Assuming (33) and (34), we have
+$\frac{K_{t}-{\mathbb{E}}(K_{t})}{\sqrt{{\rm Var}(K_{t})}}=\frac{X_{t}-{\mathbb{E}}(X_{t})}{\sqrt{{\rm Var}(X_{t})}}(1+o(1))+\frac{\sum_{s=1}^{t-1}(X_{s}-{\mathbb{E}}(X_{s}))}{\sqrt{{\rm Var}(K_{t})}}.$ (35)
+Suppose that
+$\frac{K_{t}-{\mathbb{E}}(K_{t})}{\sqrt{{\rm Var}(K_{t})}}\stackrel{{\scriptstyle d}}{{\to}}{\rm N}(0,1).$
+By our assumptions, the variance of the last term of (35) is $o(1)$ and so by Slutsky’s theorem [durrett]
+$\lim_{t\to\infty}{\mathbb{P}}\left(\frac{X_{t}-{\mathbb{E}}(X_{t})}{\sqrt{{\rm Var}(K_{t})}}\leq 0\right)=\lim_{t\to\infty}{\mathbb{P}}\left(\frac{K_{t}-{\mathbb{E}}(K_{t})}{\sqrt{{\rm Var}(X_{t})}}\leq 0\right)=1/2,$ (36)
+We will show that (36) is impossible. The c.d.f. of $X_{t}$ is given by
+$P(X_{t}\leq x_{t})$ $=$ $\sum_{k_{t}=0}^{x_{t}}{\mathbb{P}}(X_{t}=k_{t})$
+$=$ $\frac{1}{\binom{N_{t}}{n_{t}}}\sum_{k_{t}=0}^{x_{t}}\binom{N_{t-1}-1+n_{t}-k_{t}}{n_{t}-k_{t}}$
+$=$ $\frac{1}{\binom{N_{t}}{n_{t}}}\left(\sum_{j=0}^{n_{t}}\binom{N_{t-1}-1+j}{j}-\sum_{j=0}^{n_{t}-x_{t}-1}\binom{N_{t-1}-1+j}{j}\right)$
+$=$ $\frac{1}{\binom{N_{t}}{n_{t}}}\left(\binom{N_{t}}{n_{t}}-\binom{N_{t}-x_{t}-1}{n_{t}-x_{t}-1}\right)$
+$=$ $1-\frac{(N_{t}-x_{t}-1)!/(n_{t}-x_{t}-1)!}{N_{t}!/n_{t}!}$
+$\geq$ $1-\left(\frac{N_{t}-x_{t}-1}{N_{t}}\right)^{N_{t-1}},$ (39)
+where we have used (22) at (4), (17) at (4), and $(N_{t}-x_{t}-1-i)/(N_{t}-i)\leq(N_{t}-x_{t}-1)/N_{t}$ for all $0\leq i\leq N_{t-1}-1$ at (39). We therefore have
+$\liminf_{t\to\infty}P(X_{t}\leq{\mathbb{E}}(X_{t}))$ $\geq$ $1-\lim_{t\to\infty}\left(\frac{N_{t}-{\mathbb{E}}(X_{t})-1}{N_{t}}\right)^{N_{t-1}}$
+$=$ $1-\lim_{t\to\infty}\left(\frac{N_{t}-n_{t}(1+o(1)/N_{t-1}-1}{N_{t}}\right)^{N_{t-1}}$
+$=$ $1-\lim_{t\to\infty}\left(1-(1+o(1))/N_{t-1}\right)^{N_{t-1}}$
+$=$ $1-e^{-1}>1/2,$
+contradicting (36).
 5 Discussion
@@
 $\asymp$ $\frac{1}{\log n};$
-see [2].
-By considering the process of cycle counts $(C_{1}(n),C_{2}(n),\ldots,C_{n}(n))$, where $C_{i}(n)$ is the number of cycles of size $i$, we may write $K_{n}=\sum_{i=1}^{n}C_{i}(n)$, and, moreover, one may show that
-$\frac{\sum_{i=1}^{\lfloor n^{t}\rfloor}C_{i}(n)-t\log n}{\sqrt{\log n}}$
-is asymptotically standard Brownian motion; see [1].
-If suitable approximations could be made to the component count process of multiset permutations, then similar progress might be made for the results obtained in this paper when $\theta=1$.
+see [BH].
+By considering the process of cycle counts $(C_{1}(n),C_{2}(n),\ldots,C_{n}(n))$, where $C_{i}(n)$ is the number of cycles of size $i$, we may write $K_{n}=\sum_{i=1}^{n}C_{i}(n)$.
+Moreover, [DP] shows that $B_{n}(\cdot)\stackrel{{\scriptstyle d}}{{\to}}B(\cdot)$, where
+$B_{n}(t):=\frac{\sum_{i=1}^{\lfloor n^{t}\rfloor}C_{i}(n)-t\log n}{\sqrt{\log n}},\quad 0\leq t\leq 1,$
+and $B(t)$ is standard Brownian motion.
+If suitable approximations could be made to the process of cycle counts for multiset permutations, then similar progress might be made for the results obtained in this paper when $\theta=1$.
+{acknowledgement}
+[title=Acknowledgments] The author thanks the anonymous referee and the Associated Editor for a careful reading of this paper and for helpful suggestions.